OpenStudy (anonymous):

how to find the electric field at a point x above the axis of a uniformly charged semi circular ring?

OpenStudy (masumanwar):

what ever charge you can find at any finite distance using coulombs law

OpenStudy (anonymous):

i know that too but the problem in 3D becomes a bit confusing

OpenStudy (anonymous):



OpenStudy (anonymous):

dq = Q/2piR

OpenStudy (anonymous):

kdq/r^2 = kQ/2piR *{ cos(theta) = z/(sqrt(z^2+R^2)) } * 1/(z^2+R^2) integrate all around the loop: kQ/(2piR) * (z/ (z^2 +R^2)^(3/2)) ds =kQ/(2piR) * (z/ (z^2 +R^2)^(3/2)) * 2piR

OpenStudy (anonymous):

there are 3 components of EF in space ,for a full ring two gets cancelled but for a half ring i think only one gets cancelled is it true?

OpenStudy (anonymous):

ah half ring? I didn't see that part..

OpenStudy (anonymous):

can you still help?

OpenStudy (anonymous):

you'll notice that one of the coordinate components of the field still cancels...so for instance..